# Closed-Form Feedback-Free Learning with Forward Projection

### v3

@@ Results
Table 1: Test performance of various learning methods across different datasets. The FMNIST dataset was modelled with a multi-layer perception; Promoters and PTBXL-MI datasets were modelled using 1D convolutional neural networks (1D-CNNs); CIFAR2 (first two CIFAR10 classes only) was modelled using vision transformers and CIFAR10 was modelled with 2D-CNNs. BP: bacpropagation; DTP: Difference Target Propagation; $F$: Forward Forward; FP: Forward Projection; LS: Local Supervision; PC: Predictive Coding; RF: random features
Alternative Feedback-free Approaches
To assess the value of the FP target generation function, we compared the performance of closed-form regression models fitted to targets generated by alternative functions, including simple label projection ($\tilde{\textbf{Z}}_{l}:=\textbf{yU}_{l}$) and label projection with additive noise ($\tilde{\textbf{Z}}_{l}:=\textbf{yU}_{l}+\textbf{E}$). To evaluate the capability of feedback-free training methods to handle information “bottlenecks”, they were applied to optimise MLP architectures for FMNIST classification. MLPs were generated with 1000 hidden neurons in the first and second layers ($m_{1}=m_{2}=1000$), and $m_{3}\in\{100,200,400,800\}$ neurons in the final hidden layer. FP outperformed other feedback-free approaches (Figure 2-A). The performance of RF deteriorated in models with small penultimate layers, as relevant information was less likely to be represented by random projection [23]. The performance of simple label projection deteriorated in models with large penultimate layers, a result which may be attributable to rank deficiency (see remark A.5 in SI). Noisy label projection maintained steady performance but at a lower level than FP.
Figure 2: A: Comparison of feedback-free fitting methods on FMNIST. MLP architectures had 1000 neurons in the first and second layers and 100, 200, 400 or 800 neurons in the final hidden layer. B-D: Test performance of few-shot trained 2D-CNN models. Mean test AUC is reported over 50 few-shot training experiments. B: Chest X-ray (CXR) task. C: Optical Coherence Tomography (OCT) task. D: CIFAR2 task in which models were required to classify the first two classes (aeroplane and automobile). Mean test AUC is reported over 50 few-shot training experiments. Models were fitted with $N\in\{5,10,15,20,30,40,50\}$ training samples from each class in CXR and OCT tasks and $N\in\{25,50,75,100\}$ samples per class for the CIFAR2 task. Predictive Coding and Difference Target Propagation are plotted separately in Supplementary Figure 7. BP: backpropagation; $F$: Forward-Forward; FP: Forward Projection; LS: Local Supervision; RF: random features.
Few-Shot Learning
“Few-shot” learning is a constrained learning scenario in which the number of data samples available for training is small. High-dimensional data, such as images, pose a challenge for few-shot learning, as many spurious features may exist. Thus, successful few-shot training methods must select generalisable features in the presence of these confounders. Few-shot learning was assessed in image classification tasks using a 2D-CNN architecture. The optical coherence tomography task (OCT) [30] required that the model discriminate images of healthy retinas from those with choroid neovascularization, a pathology that affects the eye and manifests mainly as abnormal growth of blood vessels behind the retina. The paediatric chest X-ray (CXR) task [30] required the model to discriminate between images of viral pneumonia, bacterial pneumonia, or healthy controls. The CIFAR2 task required the model to discriminate between the first two CIFAR10 classes (aeroplane and automobile). Few-shot training datasets were generated by subsampling, with $N\in\{5,10,15,20,30,40,50\}$ training examples from each class for OCT and CXR tasks, and $N\in\{25,50,75,100\}$ training examples per class for the CIFAR2 task. Data augmentation was not employed. Model generalisability was assessed on all test data (CXR: $N_{\text{test}}=431$; OCT: $N_{\text{test}}=327$, CIFAR2: $N_{\text{test}}=2000$). Forward Projection-trained models demonstrated the greatest few-shot generalisability in CXR (Figure 2-B), OCT (Figure 2-C), outperforming all other methods, including backpropagation, in all the tasks with $N\leq 40$ training samples. In the CIFAR few-shot task, Forward Projection achieved the best performance of any local learning method in experiments with $N\in\{50,75,100\}$, being outperformed only by backpropagation. With as few as $N=10$ training samples, Forward Projection fitted discriminative models for the classification of OCT (test AUC: $84.5\pm 5.8$) and CXR (test AUC: $76.6\pm 5.8$). CXR and OCT datasets highlighted two distinct vulnerabilities of backpropagation training in few-shot conditions. In the $N=5$ setting on OCT, backpropagation overfitted the training samples (Train AUC: $86.8\pm 13.2$; Test AUC: $61.8\pm 17.4$), as models integrated noise into decision functions. On the other hand, backpropagation failed to achieve adequate model fitting in the $N=10$ setting on CXR (Train AUC: $69.0\pm 8.1$; Test AUC: $59.7\pm 6.2$). Random features performed comparably to backpropagation in tasks with fewest training samples, but performance improved only slightly with larger sample sizes. Random Features models could not overfit within convolutional layers, as these contained no free parameters. However, Random Features had limited capacity to learn structural features in larger training samples. The improvement that Forward Projection provided over Random Features is therefore attributable to structural feature learning within hidden convolutional layers. Label Projection and Noisy Label Projection failed to converge in few-shot training (training AUC $\approx 0.50$). Neither Local Supervision nor Forward-Forward training matched the baseline generalisability of Random Features in OCT and CXR tasks - however Local Supervision achieved comparable test discrimination to Forward Projection in the CIFAR2 task when $N\geq 75$. Predictive coding and Difference Target Propagation were yielded uninformative models in few-shot learning tasks (Figure 7). Few-shot performance is tabulated in Supplemental Table 4.
Feature Interpretability
Explainability is a central issue with backpropagation-based learning, as relationships between hidden activations and model predictions may be non-monotonic, complicating the interpretation of hidden neural activities [31]. An important advantage of Forward Projection is the interpretability of hidden neuron activity with respect to label predictions (see remark A.7 in SI). Assuming $\textbf{z}_{l}\approx\tilde{\textbf{z}_{l}}$, neural potentials may be interpreted as a local label prediction $\hat{\textbf{y}}_{l}$ (Figure 1-D), such that
$\hat{\textbf{y}}_{l}:=g_{l}^{-1}(\tilde{\textbf{z}}_{l}-g_{l}(\textbf{a}_{l-1}\textbf{Q}_{l}))\textbf{U}_{l}^{+}.$ (3)
Here, $\textbf{U}_{l}^{+}$ is the Moore-Penrose generalised inverse of the label projection matrix. Likewise, pre-synaptic inputs are encoded in neural pre-activation potentials, with an analogous reconstruction function (see remark A.7 in SI). It is noted that (3) may be uninformative if $\frac{\lVert\textbf{z}_{l}-\tilde{\textbf{z}}_{l}\rVert}{\lVert\tilde{\textbf{z}}_{l}\rVert}$ is large – i.e. if $\textbf{W}_{l}$ did not achieve a good fit. Measurement of local error in training data may provide insights into the reliability of (3) during inference. Interpretation of pre-activation potentials ${z}_{l}$ is simplified by selection of bijective functions for $g_{l}$ so that $g_{l}^{-1}$ exists everywhere. In practice, a surrogate approximation to the functional inverse was observed to suffice in our experiments; for example, $g_{l}^{-1}(\cdot)\approx\operatorname{tanh}(\cdot)$ was employed as a surrogate inverse for $g_{l}(\cdot)=\operatorname{sign}(\cdot)$. In our experiments, hidden neurons of models fitted with Forward Projection were interpretable as label predictions. In the FMNIST task, test accuracy of layer explanations was observed to improve between early layers and subsequent layers (Figure 6-A), demonstrating synergistic learning. In the PTBXL-MI task, applying the surrogate layer explanation function (described in equation(3)) to the convolutional layers identified various clinically salient features for diagnosing myocardial infarction (MI). MI, a clinical condition characterised by damage to heart muscles due to poor blood flow, may manifest in ECG data with various electro-physical abnormalities. Consequently, the model must learn several distinct pathological waveform features, including elevation of the so-called “ST” segment, or inversion of the “T”-wave. Figure 3-A shows the layer explanation functions as a function of time in four patients, three of whom were diagnosed with myocardial infarction. Patient A demonstrates ST-segment depression in lead II, which is temporally consistent with peaks in the model explanation functions at each layer. Likewise, the model explanation functions peak during ST-segment elevation in patient B and during QRS widening and T-wave inversion in Patient C. In contrast, the model explanation function is near zero in Patient D, who had normal ECG morphology. Forward Projection model explanations derived using (3) from the sixth convolutional layer ($\hat{y}_{6}$) were compared with GradCAM outputs of backpropagation-trained networks. The first 15 MI-positive test instances in the PTBXL dataset were annotated by a medical doctor to segment the diagnostically relevant subsequences. Layer explanations derived by Forward Projection achieved similar discrimination performance (AUC: $0.61\pm 0.14$, AUPR: $0.52\pm 0.19$) to GradCAM (AUC: $0.59\pm 0.12$, AUPR: $0.55\pm 1.3$) (Figure 3-B).
Figure 3: A: Visualisation of layer explanations over time in a 1D-convolutional neural network trained by Forward Projection to detect myocardial infarction (MI) in electrocardiograms (ECGs) from PTBXL data. Patients A, B, C, E (diagnosed MI) and patient D (no disease) were extracted from test data. Explanations were extracted from the second, fourth and sixth convolutional layers ($\hat{y}_{2},\hat{y}_{4},\hat{y}_{6}$) using equation (3). Explanations increase with MI features (highlighted in red), including ST-segment depression (Patient A), ST-segment elevation (Patient B) and QRS widening with T-wave inversion (Patient C). B: Comparison of Forward Projection with GradCAM. Above: ECG data from Patient E (diagnosed MI), showing ST-segment elevation. Middle: Sixth convolutional layer explanation ($\hat{y}_{6}$) from a model trained by Forward Projection. Below: GradCAM output for the sixth convolutional layer of a model trained by backpropagation.
Choroid neovascularization (CNV) is the growth of abnormal blood vessels behind the retina due to diseases such as age-related macular degeneration [32]. In OCT images, CNV may be represented by various image features, including hyper-reflective dots and detachment of the retinal pigment epithelium [33]. In the OCT task, the model’s explanation functions identified regions of interest related to CNV (Figure 4). 2D-CNN models trained with only 100 instances per class learned to localise fine-grained CNV features including retinal/subretinal fluid (Patients A-C), hard exudates (Patient B), and fibrosis (Patient C).
Figure 4: Visualisation of layer explanations over space in 2D-CNNs trained by Forward Projection to detect choroid neovascularization (CNV) in the OCT task. Ensemble average of five models shown. Patients A-C, (diagnosed CNV) and patient D (no disease) were extracted from test data. Explanations were extracted from the second, fourth and sixth convolutional layers ($\hat{y}_{2},\hat{y}_{4},\hat{y}_{6}$), using (3). CNV heat-maps demonstrate high values (red) over CNV features, including retinal/sub-retinal fluid (Patients A-C) and hard exudates (Patient B), and fibrosis (Patient C), with low values (blue) over healthy retina (Patient D).
Training Complexity
We now analyse the complexity of training with the Forward Projection algorithm for a classification task by estimating the storage and computational requirements for a densely connected $m\times m$ hidden layer. We consider a dataset comprising of $N$ training samples, with label dimension $m_{L}$. Note that FP model weights can be obtained in one pass over the $N$ training samples, while all other methods require each training sample to be fed to the network for several epochs (denoted $N_{e}$). The training procedure for each layer is presented in Supplementary Figure 5.
The memory requirement for Forward Projection is $\mathcal{O}(m^{2})$ for the layer weights, $\mathcal{O}(m^{2})$ for the $Q$ matrix, and $\mathcal{O}(mm_{L})$ for the $U$ matrix. As the $\textbf{A}_{l-1}^{T}\textbf{A}_{l-1}$ and $\textbf{A}_{l-1}^{T}\tilde{\textbf{Z}}_{l}$ terms in (2) can be accumulated sequentially over data batches (See Remark A.6 in SI), two $m\times m$ matrices suffice for their storage, thereby avoiding storage of the $N\times m$ matrices $\textbf{A}_{l-1}$ and $\tilde{\textbf{Z}}_{l}$. As with all methods, Forward Projection requires $\mathcal{O}(m^{2})$ multiply-and-accumulate (MAC) operations to calculate the activations for the downstream layer in the forward pass. To generate target potentials during the forward pass, Forward Projection also requires $\mathcal{O}(m^{2})$ additional MAC operations to project through $Q$ and $\mathcal{O}(mm_{L})$ MAC operations to project through U. Computation of the $\textbf{A}_{l-1}^{T}\textbf{A}_{l-1}$ and $\textbf{A}_{l-1}^{T}\tilde{\textbf{Z}}_{l}$ terms require $\mathcal{O}(Nm^{2})$ MAC operations each. After all samples have been observed, computation of model weights using equation (2) requires $\mathcal{O}(m^{3})$ operations to invert the $\textbf{A}_{l-1}^{T}\textbf{A}_{l-1}$ term and $\mathcal{O}(m^{3})$ operations for matrix multiplication to complete the regression; however, this happens only once for each layer. Furthermore, Forward Projection requires no backward pass.
For backpropagation, $\mathcal{O}(m^{2})$ memory is required for storing layer weights, $\mathcal{O}(m^{2})$ for accumulated gradients, and $\mathcal{O}(m)$ for activations. Here, the computation requirements scale as $\mathcal{O}(N_{e}m^{2})$ MACs each for the forward pass, backward pass, and weight update calculations. Hence, in the typical setting where $m>>m_{L}$, we note that Forward Projection and backpropagation have similar memory requirements, scaling as $\mathcal{O}(m^{2})$. Notably, compute scales as $\mathcal{O}(m^{3})$ for FP versus $\mathcal{O}(N_{e}m^{2})$ for BP. For example, training a dense hidden layer with $1000$ inputs and $1000$ outputs on the FMNIST dataset ($N=60,000,m_{L}=10$) by Forward Projection requires a total of $1.2\times 10^{11}$ MAC operations for the forward pass and $1.2\times 10^{11}$ operations for the weight update. Training the same layer via 100 epochs of backpropagation requires $6.0\times 10^{12}$ MAC operations for forward passes and $1.8\times 10^{13}$ operations for weight updates. Forward projection requires only a single training epoch, leading to a proportional speed up in the overall computation time. Note that Local Supervision and Forward-Forward algorithms have similar computational and memory requirements as backpropagation, as illustrated in Table 2. We also present the wall-clock time for full training of an MLP with $3\times 1000$ hidden neurons on FMNIST, demonstrating a $66\times$ speedup with Forward Projection in this example. Further timing results are provided in Table5 in SI. Computations were run using the Google Colab service with an NVIDIA T4 graphics processing unit. Training complexities of other methods are discussed in remark A.12.
@@
Discussion
Few-shot learning is ubiquitous in biological systems, which exhibit rapid neuronal adaptation to changing environments [35]. In our experiments, Forward Projection demonstrated clear performance advantages over backpropagation and local learning approaches in few-shot learning tasks, presenting a plausible method for learning new tasks rapidly. In this setting, convolutional features learned by Forward Projection yielded more generalisable models than backpropagation, which overfitted in some experiments and underfitted in others, even underperforming random features in some cases. Forward Projection performed similarly to Local Supervision and Forward-Forward training in these tasks. Forward Projection also maintained reasonable performance on activation functions which were untrainable by SGD-based methods (see section A.13 in SI).
The closed-form Forward Projection fit is computable in a single pass over the data for each layer, presenting an opportunity to expedite training and reduce environmental footprint. The efficiency of FP training is attributable to a substantially different operational sequence to that employed in SGD-based approaches. Firstly, FP training collects a gram matrix of pre-synaptic activity over a single epoch, which is subsequently inverted during a one-step weight matrix computation. Secondly, FP completes fitting for each layer before initialising successive layers. In contrast, SGD-based approaches fit all the layers in the network iteratively, using feedback from neuronal outputs and downstream layers. FP training requires no retrograde communication between neural output activations and pre-synaptic connections, enabling direct training on hardware with unidirectional synaptic and neuronal communication. This differs from current iterative local learning methods, which require backward communication from neuronal outputs to optimise pre-synaptic parameters (see remark A.1 in SI). Although FP addresses the feedback-free constraint inherent in biological learning systems, the biological plausibility of one-step learning via matrix inversion remains uncertain. Iterative fitting via gradient descent in sequential, batch-based learning offers a practical and scalable approach for aligning weights with locally generated targets, as defined in (4) (see remark A.10 in SI). Although closed-form optimisation remains a challenge for recurrent architectures, future research will explore integrating iterative training with random temporal convolutions to generalise the FP framework to dynamic models and expand its applicability to temporal learning tasks. This backpropagation-free strategy holds promise for biologically inspired computing systems, such as spiking neural networks, where non-differentiable activation functions preclude standard backpropagation [36, 37, 38]. FP training may be applicable as a pre-training step to reduce the number of training epochs required for backpropagation. In conclusion, FP is an efficient approach for neural network optimisation, employing techniques from randomised projective embedding and linear regression to fit weight matrices in one epoch using a single-step solution. Interpretability of hidden neurons in FP-trained models may be employed to improve the explainability of neural network predictions.
Methods
PTBXL-MI
The PTB-XL dataset contains 12-lead electrocardiography (ECG) recordings from 18,889 participants [28, 40]. ECG recordings of ten-second duration and 100Hz sample rate were used in our experiments. The predictive task was to discriminate ECG recordings with normal waveform morphology ($N_{\text{train}}=6,451;N_{\text{test}}=721$) from those diagnosed as myocardial infarction ($N_{\text{train}}=2,707;N_{\text{test}}=268$) by a cardiologist. Data instances with uncertain diagnoses were excluded from this analysis. Following recommendations of the data providers who provided predefined participant-disjoint dataset splits, the tenth fold was held out for model testing [28]. The first 15 MI-positive ECGs in the test set were annotated by a medical doctor to identify diagnostically relevant sections for quantitative explainability evaluation.
Promoters
CXR
The paediatric pneumonia chest X-ray dataset (“CXR”) is a retrospective cohort of patients aged between one and five years, recorded in Guangzhou Women and Children’s Medical Center, Guangzhou, originally published in [30]. Chest X-ray images were recorded as part of routine care during diagnostic workup for suspected lower respiratory tract infection. During data collection, clinicians screened the images for quality and excluded those with severe artefact or corruption. Images were annotated by two expert physicians. Local institutional review board approvals were obtained. Images were loaded in greyscale, rescaled to the $[0,1]$ intensity range and resized to $128\times 128$ pixels by bilinear interpolation. Models were required to classify images “normal” ($N_{\text{train}}=1,349;N_{\text{test}}=234$), “viral pneumonia” ($N_{\text{train}}=2,538;N_{\text{test}}=242$), or “bacterial pneumonia” ($N_{\text{train}}=1,345;N_{\text{test}}=148$).
OCT
The optical coherence tomography dataset (“OCT”) is a retrospective cohort of adult patients from five ophthalmology institutions in the USA and China recorded between 2013 and 2017 during routine care, originally published in [30]. Images were initially annotated by local medical students, who had received OCT interpretation training. Subsequent annotation was performed by four ophthalmologists and two independent retinal specialists. Horizontal foveal cut images were available in portable network graphics image format. Images were loaded in greyscale, rescaled to the $[0,1]$ intensity range, and resized to $128\times 128$ pixels by bilinear interpolation. Models were required to classify images as one of “normal” ($N_{\text{train}}=2,926;N_{\text{test}}=149$) or “choroid neovascularisation” ($N_{\text{train}}=791;N_{\text{test}}=178$).
Model Training
The FMNIST dataset was modelled using an MLP with $3\times 1000$ hidden ReLU-activated neurons. Sequential datasets (PTBXL-MI and Promoters) were modelled by a 1D-CNN architecture of four convolutional blocks. Each convolutional block included two convolutional layers with kernel dimension $3$, and strides of 1 and 2 respectively. Convolutional layers in the $l$-th block had $32\times 2^{l-1}$ filters. Convolutional implementation is detailed in Supplementary section A.2. Convolutional outputs were aggregated by global average pooling in the penultimate layer. For gradient-descent-based learning algorithms, batch normalisation layers were included between convolutional blocks. CIFAR2 modelling in Table 1 used a vision transformer architecture [41] operating on image patches of dimension $4\times 4$, with a sequential stack of four multi-headed attention layers, each having 8 heads, embedding dimension 64, and MLP dimension 64. 2D-CNN architectures had kernel dimension $3\times 3$and the $l$-th convolutional block had $16\times 2^{l-1}$ filters (CXR and OCT) $32\times 2^{l-1}$ filters (CIFAR2) or $64\times 2^{l-1}$ filters (CIFAR10). The $\operatorname{sign}$ function was employed to generate target activations for Forward Projection models, such that
$\tilde{\textbf{z}}_{l}=\operatorname{sign}(\textbf{a}_{l-1}\textbf{Q}_{l})+\operatorname{sign}(\textbf{yU}_{l}).$ (4)
Data augmentation was not performed in our experiments. Models were fitted using the PyTorch library. For SGD-based methods, early stopping was performed according to validation loss. Model weight initialisation was random, and optimisation was performed with the Adam optimiser with a learning rate of 0.001, training to minimise validation loss by early stopping with a patience of five epochs. In few-shot experiments, the learning rate was reduced to 0.0001 for SGD-based models and a patience of ten epochs was employed. It is acknowledged that direct comparison of gradient descent based training with closed-form solutions is not strictly “like-for-like”. However, all efforts were made to ensure experimental conditions were otherwise equivalent. Models were fitted to minimise categorical or binary cross-entropy as appropriate. Implementations of Forward-Forward, Local Supervision, Difference Target Propagation and Predictive Coding methods are detailed in supplementary section A.4.
Explainability Analysis
Model architectures for explainability analysis were equivalent to those used for the main experiments. Pre-activation potentials in hidden neurons were interpreted as local label predictions using the approximation to (3) as
$\hat{\textbf{y}}_{l}:=\tanh(\textbf{z}_{l}-\operatorname{sign}(\textbf{a}_{l-1}\textbf{Q}_{l}))\textbf{U}_{l}^{+}.$ (5)
Declarations
Funding
This research was supported in part by the EPSRC Open Fellowship EP/X011356/1.
Competing Interests
Ethics approval and consent to participate
Data Availability
The PTBXL-MI dataset is available from PhysioNet [40]. The Promoters dataset is available from the GenomicBenchmarks library [27] and the original publication [26]. The CXR and OCT datasets are provided in [30]. Pre-processed datasets and cross-validation folds are available for download at fp_datasets - Mendeley Data [42].
Code Availability
All code required to reproduce study findings is available from github.com/robertoshea/forward_projection.
Author Contribution

### v4

@@ Acknowledgements
Table 1: Test performance of various learning methods across different datasets. The FMNIST dataset was modelled with a multi-layer perceptron; Promoters and PTBXL-MI datasets were modelled using 1D convolutional neural networks (1D-CNNs); CIFAR2 (first two CIFAR10 classes only) was modelled using vision transformers and CIFAR10 was modelled with 2D-CNNs. BP: backpropagation; DTP: Difference Target Propagation; FF: Forward-Forward; FP: Forward Projection; LS: Local Supervision; PC: Predictive Coding; RF: Random Features.
@@
Figure Legends
Figure 1. Graphical overview of forward projection. A: Forward projection algorithm for fitting layer weights $\mathbf{W}_{1},\ldots,\mathbf{W}_{l}$ to model labels $\mathbf{Y}$ from data $\mathbf{X}$. B: Procedure for generating the $l$-th layer target potentials $\tilde{\mathbf{Z}}_{l}$. Pre-synaptic inputs $\mathbf{A}_{l-1}$ and labels $\mathbf{Y}$ are projected with fixed matrices $\mathbf{Q}_{l}$ and $\mathbf{U}_{l}$, respectively, before applying non-linearity $g_{l}$. C: Optimising $\mathbf{W}_{l}$ to predict $\tilde{\mathbf{Z}}_{l}$ from $\mathbf{A}_{l-1}$ by ridge regression with penalty $\lambda$. D: Interpreting membrane potentials $\mathbf{z}_{l}$ as a local label prediction $\hat{\mathbf{y}}_{l}$ given pre-synaptic inputs $\mathbf{a}_{l-1}$ and projection matrices $\mathbf{Q}_{l}$ and $\mathbf{U}_{l}^{+}$, where $\mathbf{U}_{l}^{+}$ is the pseudo-inverse of $\mathbf{U}_{l}$.
Figure 2. Performance of Forward Projection with backpropagation and local learning approaches. A: Comparison of feedback-free fitting methods on FMNIST. MLP architectures had 1000 neurons in the first and second layers and 100, 200, 400 or 800 neurons in the final hidden layer. B-D: Test performance of few-shot trained 2D-CNN models. Mean test AUC is reported over 50 few-shot training experiments. B: Chest X-ray (CXR) task. C: Optical Coherence Tomography (OCT) task. D: CIFAR2 task in which models were required to classify the first two classes (aeroplane and automobile). Models were fitted with $N\in\{5,10,15,20,30,40,50\}$ training samples from each class in CXR and OCT tasks and $N\in\{25,50,75,100\}$ samples per class for the CIFAR2 task. Predictive Coding and Difference Target Propagation are plotted separately in Supplementary Figure 7. BP: backpropagation; FF: Forward-Forward; FP: Forward Projection; LS: Local Supervision; RF: Random Features.
Figure 3. Forward Projection layer interpretions for electrocardiogram analysis. A: Visualisation of layer explanations over time in a 1D-convolutional neural network trained by Forward Projection to detect myocardial infarction (MI) in electrocardiograms (ECGs) from PTBXL data. Patients A, B, C, E (diagnosed MI) and patient D (no disease) were extracted from test data. Explanations were extracted from the second, fourth and sixth convolutional layers ($\hat{y}_{2},\hat{y}_{4},\hat{y}_{6}$) using equation (3). Explanations increase with MI features (highlighted in red), including ST-segment depression (Patient A), ST-segment elevation (Patient B) and QRS widening with T-wave inversion (Patient C). B: Comparison of Forward Projection with GradCAM. Top: ECG data from Patient E (diagnosed MI), showing ST-segment elevation. Middle: Sixth convolutional layer explanation ($\hat{y}_{6}$) from a model trained by Forward Projection. Below: GradCAM output for the sixth convolutional layer of a model trained by backpropagation.
Figure 4. Visualisation of layer explanations over space in 2D-CNNs trained by Forward Projection to detect choroid neovascularization (CNV) in the OCT task. Ensemble average of five models shown. Patients A-C (diagnosed CNV) and patient D (no disease) were extracted from test data. Explanations were extracted from the second, fourth and sixth convolutional layers ($\hat{y}_{2},\hat{y}_{4},\hat{y}_{6}$), using (3). CNV heat-maps demonstrate high values (red) over CNV features, including retinal/subretinal fluid (Patients A-C) and hard exudates (Patient B), and fibrosis (Patient C), with low values (blue) over healthy retina (Patient D).